# Patterns of active dipolar particles in external magnetic fields

### v2

@@ III.2 Influence of an external field
Figure 4 shows typical system configurations for three different values of the magnetic moments $\mu=$ 1.0 , 1.4 and 2.61.01.42.61.01.42.6start_ARG 1.0 end_ARG , start_ARG 1.4 end_ARG and start_ARG 2.6 end_ARG (columns from left to right) at these three field strengths ($B=$ 0.1 , 1.0 and 140.11.0140.11.014start_ARG 0.1 end_ARG , start_ARG 1.0 end_ARG and start_ARG 14 end_ARG in the top, middle, bottom row, respectively) for particles with moderate self-propulsion speed ($v_{0}=23$). The orientation of the particles is indicated by their color according to the color wheel. The magnetic field points from left to right, so alignment with the field can be seen by the cyan color of the particles. Systems with a weak external magnetic field (top row) display similar configurations to what was observed in systems without external magnetic field (compare with fig. 3). In systems with intermediate external magnetic fields (middle row), particles align with the direction of the external field (as seen by their cyan color). In addition, particles assemble into chains for strong external magnetic fields ($B=14$). These chains often have defects such as branches or small loops. Here, the average chain length is seen to increase with the magnetic moment of the particles, spanning the whole system size for strong magnetic interactions and strong external field strengths (for $\mu=2.6,\,B=14$, bottom right). These columnar ordered clusters are qualitatively similar to structures formed by non-equilibrium ferrofluids with external magnetic fields [23]. In contrast to systems without an external magnetic field, these columnar clusters mostly stay evenly separated and do not combine into connected networks of chains or into broad bands (at the low density simulated here, broad bands are seen for higher density, as discussed below). In fact, we observed in systems with strong magnetic interactions ($\mu=2.6$) that an increase of the external magnetic field strength causes broad bands to separate into thinner bands with fewer lanes or into individual chains of dipolar particles (compare the cases for $\mu=2.6$, $B=1.0$ and $\mu=2.6$, $B=14$ in fig. 4).
@@
Figure 6: Diagrams of states for three different external magnetic field strengths $B=$ 0.1 , 1.0 and 140.11.0140.11.014start_ARG 0.1 end_ARG , start_ARG 1.0 end_ARG and start_ARG 14 end_ARG (left to right). Classification of states is based on combinations of order parameter criteria (see table 1). System at low density ($\Phi=0.13$).
@@
Using again the criteria based on the three order parameters, we classified the observed states of the system and summarized the results in a diagram of states, shown in fig. 6. The figure shows the diagrams of states for the three different magnetic field strengths $B=$ 0.1 , 1.0 and 140.11.0140.11.014start_ARG 0.1 end_ARG , start_ARG 1.0 end_ARG and start_ARG 14 end_ARG (left to right) and for the density $\Phi=0.13$. Based on these diagrams we can make three key observations: First, disordered gas states (gray circles) are strongly suppressed with increasing magnetic field strength. In addition, the onset of the oriented gas state (purple plus) shifts to lower values of $\mu$. Second, intermediate as well as strong magnetic fields (center and right diagram) suppress the formation of networks of chains (yellow square) and instead promote the formation of oriented states, specifically the oriented gas (purple plus), oriented chains (green pentagon) and bands (red cross, for intermediate field strength). Third, and most importantly, strong external magnetic fields suppress the formation of bands. For strong external magnetic fields (right diagram), two states dominate the diagram, the oriented gas (purple plus) and oriented chains (green pentagon). The line separating these two states resembles the corresponding line separating the gas and chain state in small confined systems of dipolar active particles investigated in previous work [15].
@@
For strong magnetic fields, band patterns were largely absent (right panel in fig. 6). We only found bands for two parameter combinations ($\mu=2.6$, $v_{0}=200$ and $\mu$=4, $v_{0}=500$, indicated by the red crosses in fig. 6). In these cases we observed the following noteworthy behavior, shown by the snapshots in fig. 7. Dipolar active particles that have self-assembled into long chains transiently show oscillations. The amplitude of these oscillations was seen to grow towards their tail of a chain. We suspect that such oscillations might be caused by a buckling instability: In the chains, particles with strong dipolar interactions are packed rather closely. The combination of high swimming speeds and translational noise can then cause a temporary compression and stress along the chain that is resolved by buckling of the chain. When the dipolar interactions were reduced, long chains were seen to break into smaller segments. These breaks typically occurred near the tail of the chain, where the oscillations reach their maximum amplitude. This observation suggests that fragmentation induced by buckling may provide a mechanism for the coexistence of chains of different lengths. We note that the breaking of chains resembles the fragmentation of passive chains observed in magnetorheological suspensions driven by a rotating field [36, 37]. In our simulations these oscillating chains are sees as a transient state towards band formation. At some point, individual oscillating chains meet. Then, alignment with the neighboring particles suppresses oscillations and a band starts to form. Again, similar lateral aggregation of individual chains into bands and columnar lanes has been observed in magnetorheological fluids [22, 23, 24, 38].
@@
Figure 7: Series of snapshots showing oscillating chains for dipolar active particles with strong magnetic interactions $\mu=4$ and high self-propulsion speed $v_{0}=500$. Snapshots taken at times 26.46 , 26.56 and 26.6626.4626.5626.6626.4626.5626.66start_ARG 26.46 end_ARG , start_ARG 26.56 end_ARG and start_ARG 26.66 end_ARG (left to right), color wheel indicates orientations. External magnetic field points from left to right, magnetic field strength was set to $B=14$.
@@ III.3 Effect of density
Figure 8: Diagrams of states for three different external magnetic field strengths $B=$ 0.1 , 1.0 and 140.11.0140.11.014start_ARG 0.1 end_ARG , start_ARG 1.0 end_ARG and start_ARG 14 end_ARG (left to right). Classification of states is based on combinations of order parameter criteria (see table 1). System at intermediate density ($\Phi=0.23$). Cut-off for cluster formation set at $\delta_{c}=1.21$.
For systems of high density (our case with $\Phi=0.57$), three phenomena compete with one another: polar order, chain formation and MIPS, as pointed out by Liao et al. [14]. The addition of an external magnetic field introduces global polar order and suppresses formation of complex networks of chains. Instead, we observed that strong external magnetic field strengths of $B=14$ clearly organize the system with high magnetic moments into columnar structures (see fig. 4 $\mu=2.6,B=14$). Because of the increase in density, these columnar structures are made up of multiple lanes in off-register alignment and not of single lanes, as seen in systems at low density.
In the diagram of states, we observed that the increased density causes the previously suppressed band states (red crosses) to reappear. In fact, bands dominate the corresponding diagram of states for strong external magnetic fields and are independent of activity (fig. 9, right). Off-register band formation has been previously observed in passive dipolar systems with constant external magnetic fields [39, 40, 41]. In systems at high density, a clear discrimination between bands and columnar states can no longer be made. A limiting factor might be the finite system size, as observed in other systems of active particles [42].
Figure 9: Diagrams of states for three different external magnetic field strengths $B=$ 0.1 , 1.0 and 140.11.0140.11.014start_ARG 0.1 end_ARG , start_ARG 1.0 end_ARG and start_ARG 14 end_ARG (left to right). Classification of states is based on combinations of order parameter criteria (see table 1). System at high density ($\Phi=0.57$). Cut-off for cluster formation set at $\delta_{c}=1.05$.
@@ III.4 Columnar clusters
Figure 10: Snapshots of system with intermediate density and strong magnetic interactions show columnar clusters, parallel bands spanning the whole system in the direction of the field that shrink with increasing external field. The density is $\Phi=0.23$, the velocity $v=23$.
Snapshots of the simulations indicate (see fig. 10) that the number of lanes, i.e., of parallel chains in each cluster can vary and depends on the field strength. The clusters appear to be quite regularly spaced in the direction perpendicular to the external magnetic field. All these properties are reminiscent of magnetic field-induced columnar clusters in ferrofluids [22, 23, 24].
@@
Figure 12: Columnar spacing $\overline{\Delta r_{y}}$ (left) and average number of lanes per cluster $\overline{n_{\mathrm{l}}}$ (right) against external magnetic field strength $B$. The magnetic moment was set to $\mu=4$ and density to $\Phi=0.23$. Colors indicate increasing self-propulsion speed of the particles. Data sampled over trajectories at times $t=18-20$ every $\Delta t=0.1$. Data points show systems that fulfill criteria $\left\langle\phi_{e}\right\rangle>0.5$ and $\left\langle\phi_{p}\right\rangle>0.5$.
Figure 12 shows the columnar spacing $\overline{\Delta r_{y}}$ and the average number of lanes per column $\overline{n_{\mathrm{l}}}$ for systems with strong magnetic interactions $\mu=4$. Both the columnar spacing and the number of lanes per cluster decay rapidly and approach a fixed value as the external magnetic field strength increases. Activity seems to have only a minor effect on this behavior. Columnar spacing has been reported to be independent of the external magnetic field in experiments with (passive) ferrofluids [24]. Our result for active particles at for strong magnetic fields mirror this behavior. However, we observed a strong dependence of the columnar spacing in the regime of weak external magnetic fields. We note that in some cases, a weak external magnetic field seems to promote separation between columnar structures.
As reported in [24], columnar spacing is reduced with increasing density in ferrofluids. While we did qualitatively observe in movies of our simulations that the number of lanes in each column increases with density, more systematic investigation of density effects would be needed to make quantitative statements about the influence of density on columnar structures.
@@ IV Conclusions
In this study, we used extensive simulations to investigate the effect of a homogeneous external magnetic field on systems of active dipolar particles. Based on the combination of three order parameters, we determined diagrams of state for different magnetic field strengths to provide a comprehensive overview of the collective behaviors observed in such system. Consistent with previous work [14, 15], we see that dipolar interactions without an external field result in chains, which at higher densities form networks and bands. Motility-induces phase separation is only observed in the absence of dipolar interactions, already the weakest interaction we simulated resulted in networks of chains instead. The dipolar interaction likely prevents the jammed configurations and, thus, the slow-down required for motility-induced phase separation.
@@
Under the influence of an external magnetic field oriented chains are dominant and bands are seen less frequently. For weak interactions, an oriented gas of individual dipoles is observed. Both states are only weakly dependent on the the level of activity and practically independent of it when a strong external magnetic field is present. This observation indicates that for strong field, the magnetic interactions dominate over activity, such that systems of active dipolar particles resemble passive ferrofluids more than classical (non-dipolar) active Brownian particles. This observation for strong field stands in stark contrast to the chain formation observed in systems without external magnetic fields, where the activity plays a crucial role in collective phenomena [14, 15, 18].
@@
Two observations may merit further research in future work: (i) We observed oscillating chains as a transient state towards band formation in our system. We suspect that these oscillations are caused by a buckling instability and that fragmentation of long chains due to buckling allows chains of different lengths to coexist in dilute systems, despite high magnetic moments. The self-assembled active dipolar chains we study here have some similarities to active polymers, in which the monomers are linked by a spring potential [43]. In that case, buckling has been studied both theoretically and experimentally [44, 45, 46, 47] and was shown to be controlled by the flexure number, which characterizes the relation between active propulsion and bending rigidity [44, 45].
@@
(ii) We observed that active dipolar particles form columnar clusters that resemble structures induced by external magnetic fields in (passive) ferrofluids [22, 23, 24]. A quantitative analysis showed that the number of lanes per column and the columnar spacing decreases rapidly with increasing magnetic field strength, but become constant for strong fields. The strong-field behavior is similar to observations in ferrofluids, where Columnar spacing has been reported to be independent of the external magnetic field strength [22, 24]. Further features of the system and experimental protocols (such as the application of the external magnetic field, presence of hydrodynamic interactions, polydispersity and boundary conditions) may have an effect on this result. Here, in a first step, we have shown here activity does not strongly affect separation between columns. In future studies, the effect of density could be investigated in more detail and compared to observations for ferrofluids [22].
@@
Our results together with those from previous studies [14, 17] show that a preferred orientation (introduced by the external field) coupled to activity can significantly change the observed collective behaviors. In our model, activity is introduced as an active force acting in the direction of the magnetic moment of the particle. Therefore, orientation competes with the formation of structures based on random or conflicting directions of self-propulsion such as MIPS. These observations indicate that the details of how activity is introduced into the system and whether or how the direction of self-propulsion is coupled to the direction of the dipoles should be important for the type of patterns and collective phenomena that are observed. They suggest that even richer collective phenomena are possible if there is loose coupling between the dipolar and the self-propulsion direction of the active particles.

### v4

@@ III.2 Influence of an external field
Figure 4 shows typical system configurations for three different values of the magnetic moments $\mu=$ 1.0 , 1.4 and 2.61.01.42.6 (columns from left to right) at these three field strengths ($B=$ 0.1 , 1.0 and 140.11.014 in the top, middle, bottom row, respectively) for particles with moderate self-propulsion speed ($v_{0}=23$). The orientation of the particles is indicated by their color according to the color wheel. The magnetic field points from left to right, so alignment with the field can be seen by the cyan color of the particles. Systems with a weak external magnetic field (top row) display similar configurations to what was observed in systems without external magnetic field (compare with fig. 3). In systems with intermediate external magnetic fields (middle row), particles align with the direction of the external field (as seen by their cyan color). In addition, particles assemble into chains for strong external magnetic fields ($B=14$). These chains often have defects such as branches or small loops. Here, the average chain length is seen to increase with the magnetic moment of the particles, spanning the whole system size for strong magnetic interactions and strong external field strengths (for $\mu=2.6,\,B=14$, bottom right). These columnar ordered clusters are qualitatively similar to structures formed by non-equilibrium ferrofluids with external magnetic fields [Flores1999]. In contrast to systems without an external magnetic field, these columnar clusters mostly stay evenly separated and do not combine into connected networks of chains or into broad bands (at the low density simulated here, broad bands are seen for higher density, as discussed below). In fact, we observed in systems with strong magnetic interactions ($\mu=2.6$) that an increase of the external magnetic field strength causes broad bands to separate into thinner bands with fewer lanes or into individual chains of dipolar particles (compare the cases for $\mu=2.6$, $B=1.0$ and $\mu=2.6$, $B=14$ in fig. 4).
@@
Figure 6: Diagrams of states for three different external magnetic field strengths $B=$ 0.1 , 1.0 and 140.11.014 (left to right). Classification of states is based on combinations of order parameter criteria (see table 1). System at low density ($\Phi=0.13$).
@@
Using again the criteria based on the three order parameters, we classified the observed states of the system and summarized the results in a diagram of states, shown in fig. 6. The figure shows the diagrams of states for the three different magnetic field strengths $B=$ 0.1 , 1.0 and 140.11.014 (left to right) and for the density $\Phi=0.13$. Based on these diagrams we can make three key observations: First, disordered gas states (gray circles) are strongly suppressed with increasing magnetic field strength. In addition, the onset of the oriented gas state (purple plus) shifts to lower values of $\mu$. Second, intermediate as well as strong magnetic fields (center and right diagram) suppress the formation of networks of chains (yellow square) and instead promote the formation of oriented states, specifically the oriented gas (purple plus), oriented chains (green pentagon) and bands (red cross, for intermediate field strength). Third, and most importantly, strong external magnetic fields suppress the formation of bands. For strong external magnetic fields (right diagram), two states dominate the diagram, the oriented gas (purple plus) and oriented chains (green pentagon). The line separating these two states resembles the corresponding line separating the gas and chain state in small confined systems of dipolar active particles investigated in previous work [Telezki2020].
@@
For strong magnetic fields, band patterns were largely absent (right panel in fig. 6). We only found bands for two parameter combinations ($\mu=2.6$, $v_{0}=200$ and $\mu$=4, $v_{0}=500$, indicated by the red crosses in fig. 6). In these cases we observed the following noteworthy behavior, shown by the snapshots in fig. 7. Dipolar active particles that have self-assembled into long chains transiently show oscillations. The amplitude of these oscillations was seen to grow towards their tail of a chain. We suspect that such oscillations might be caused by a buckling instability: In the chains, particles with strong dipolar interactions are packed rather closely. The combination of high swimming speeds and translational noise can then cause a temporary compression and stress along the chain that is resolved by buckling of the chain. When the dipolar interactions were reduced, long chains were seen to break into smaller segments. These breaks typically occurred near the tail of the chain, where the oscillations reach their maximum amplitude. This observation suggests that fragmentation induced by buckling may provide a mechanism for the coexistence of chains of different lengths. We note that the breaking of chains resembles the fragmentation of passive chains observed in magnetorheological suspensions driven by a rotating field [Melle2003, Melle2003a]. In our simulations these oscillating chains are sees as a transient state towards band formation. At some point, individual oscillating chains meet. Then, alignment with the neighboring particles suppresses oscillations and a band starts to form. Again, similar lateral aggregation of individual chains into bands and columnar lanes has been observed in magnetorheological fluids [Liu1995, Flores1999, Ivey2001, DeVicente2011].
@@
Figure 7: Series of snapshots showing oscillating chains for dipolar active particles with strong magnetic interactions $\mu=4$ and high self-propulsion speed $v_{0}=500$. Snapshots taken at times 26.46 , 26.56 and 26.6626.4626.5626.66 (left to right), color wheel indicates orientations. External magnetic field points from left to right, magnetic field strength was set to $B=14$.
@@ III.3 Effect of density
Figure 8: Diagrams of states for three different external magnetic field strengths $B=$ 0.1 , 1.0 and 140.11.014 (left to right). Classification of states is based on combinations of order parameter criteria (see table 1). System at intermediate density ($\Phi=0.23$). Cut-off for cluster formation set at $\delta_{c}=1.21$.
For systems of high density (our case with $\Phi=0.57$), three phenomena compete with one another: polar order, chain formation and MIPS, as pointed out by Liao et al. [Liao2020]. The addition of an external magnetic field introduces global polar order and suppresses formation of complex networks of chains. Instead, we observed that strong external magnetic field strengths of $B=14$ clearly organize the system with high magnetic moments into columnar structures (see fig. 4 $\mu=2.6,B=14$). Because of the increase in density, these columnar structures are made up of multiple lanes in off-register alignment and not of single lanes, as seen in systems at low density.
In the diagram of states, we observed that the increased density causes the previously suppressed band states (red crosses) to reappear. In fact, bands dominate the corresponding diagram of states for strong external magnetic fields and are independent of activity (fig. 9, right). Off-register band formation has been previously observed in passive dipolar systems with constant external magnetic fields [Mohebi1996, Weis2003, Weis2005]. In systems at high density, a clear discrimination between bands and columnar states can no longer be made. A limiting factor might be the finite system size, as observed in other systems of active particles [Chate2008].
Figure 9: Diagrams of states for three different external magnetic field strengths $B=$ 0.1 , 1.0 and 140.11.014 (left to right). Classification of states is based on combinations of order parameter criteria (see table 1). System at high density ($\Phi=0.57$). Cut-off for cluster formation set at $\delta_{c}=1.05$.
@@ III.4 Columnar clusters
Figure 10: Snapshots of system with intermediate density and strong magnetic interactions show columnar clusters, parallel bands spanning the whole system in the direction of the field that shrink with increasing external field. The density is $\Phi=0.23$, the velocity $v_{0}=23$.
Snapshots of the simulations indicate (see fig. 10) that the number of lanes, i.e., of parallel chains in each cluster can vary and depends on the field strength. The clusters appear to be quite regularly spaced in the direction perpendicular to the external magnetic field. All these properties are reminiscent of magnetic field-induced columnar clusters in ferrofluids [Liu1995, Flores1999, Ivey2001].
@@
Figure 12: Columnar spacing $\overline{\Delta r_{y}}$ (left) and average number of lanes per cluster $\overline{n_{\mathrm{l}}}$ (right) against external magnetic field strength $B$. The magnetic moment was set to $\mu=4$ and density to $\Phi=0.23$. Colors indicate increasing self-propulsion speed of the particles. Data sampled over trajectories at times t=1820t=18-20 every $\Delta t=0.1$. Data points show systems that fulfill criteria $\left\langle\phi_{e}\right\rangle>0.5$ and $\left\langle\phi_{p}\right\rangle>0.5$.
Figure 12 shows the columnar spacing $\overline{\Delta r_{y}}$ and the average number of lanes per column $\overline{n_{\mathrm{l}}}$ for systems with strong magnetic interactions $\mu=4$. Both the columnar spacing and the number of lanes per cluster decay rapidly and approach a fixed value as the external magnetic field strength increases. Activity seems to have only a minor effect on this behavior. Columnar spacing has been reported to be independent of the external magnetic field in experiments with (passive) ferrofluids [Ivey2001]. Our result for active particles at for strong magnetic fields mirror this behavior. However, we observed a strong dependence of the columnar spacing in the regime of weak external magnetic fields. We note that in some cases, a weak external magnetic field seems to promote separation between columnar structures.
As reported in [Ivey2001], columnar spacing is reduced with increasing density in ferrofluids. While we did qualitatively observe in movies of our simulations that the number of lanes in each column increases with density, more systematic investigation of density effects would be needed to make quantitative statements about the influence of density on columnar structures.
@@ IV Conclusions
In this study, we used extensive simulations to investigate the effect of a homogeneous external magnetic field on systems of active dipolar particles. Based on the combination of three order parameters, we determined diagrams of state for different magnetic field strengths to provide a comprehensive overview of the collective behaviors observed in such system. Consistent with previous work [Liao2020, Telezki2020], we see that dipolar interactions without an external field result in chains, which at higher densities form networks and bands. In our numerical experiments, motility-induced phase separation is only observed in the absence of dipolar interactions. Adding dipolar interactions, even with the weakest interaction strength we simulated, resulted in a network of chains instead. Therefore, we conclude that dipolar interactions likely prevent the directionally disordered jamming of particles and thereby inhibit the slow-down required for motility-induced phase separation.
@@
Under the influence of an external magnetic field oriented chains are dominant and bands are seen less frequently. For weak interactions, an oriented gas of individual dipoles is observed. Both states are only weakly dependent on the the level of activity and practically independent of it when a strong external magnetic field is present. This observation indicates that for strong field, the magnetic interactions dominate over activity, such that systems of active dipolar particles resemble passive ferrofluids more than classical (non-dipolar) active Brownian particles. This observation for strong field stands in stark contrast to the chain formation observed in systems without external magnetic fields, where the activity plays a crucial role in collective phenomena [Liao2020, Telezki2020, Maloney2020].
@@
Two observations may merit further research in future work: (i) We observed oscillating chains as a transient state towards band formation in our system. We suspect that these oscillations are caused by a buckling instability and that fragmentation of long chains due to buckling allows chains of different lengths to coexist in dilute systems, despite high magnetic moments. The self-assembled active dipolar chains we study here have some similarities to active polymers, in which the monomers are linked by a spring potential [winkler2020physics]. In that case, buckling has been studied both theoretically and experimentally [Prathyusha2018, abbaspour2023effects, kurjahn2023quantifying, vliegenthart2020filamentous] and was shown to be controlled by the flexure number, which characterizes the relation between active propulsion and bending rigidity [Prathyusha2018, abbaspour2023effects].
@@
(ii) We observed that active dipolar particles form columnar clusters that resemble structures induced by external magnetic fields in (passive) ferrofluids [Liu1995, Flores1999, Ivey2001]. A quantitative analysis showed that the number of lanes per column and the columnar spacing decreases rapidly with increasing magnetic field strength, but become constant for strong fields. The strong-field behavior is similar to observations in ferrofluids, where columnar spacing has been reported to be independent of the external magnetic field strength [Liu1995, Ivey2001]. Further features of the system and experimental protocols (such as the application of the external magnetic field, presence of hydrodynamic interactions, polydispersity and boundary conditions) may have an effect on this result. In a first step, we have shown here that activity does not strongly affect separation between columns. In future studies, the effect of density could be investigated in more detail and compared to observations for ferrofluids [Liu1995].
@@
Our results together with those from previous studies [Liao2020, sese2022impact] show that a preferred orientation (introduced by the external field) coupled to activity can significantly change the observed collective behaviors. In our model, activity is introduced as an active force acting in the direction of the magnetic moment of the particle. Therefore, orientation competes with the formation of structures based on random or conflicting directions of self-propulsion such as MIPS. These observations indicate that the details of how activity is introduced into the system and whether or how the direction of self-propulsion is coupled to the direction of the dipoles should be important for the type of patterns and collective phenomena that are observed. They suggest that even richer collective phenomena are possible if there is loose coupling between the dipolar and the self-propulsion direction of the active particles.
Data and code availability
Data and code are available via GRO.data [data, code].